# Large Knowledge Model: Perspectives and Challenges

### v1

@@ Abstract
Humankind’s understanding of the world is fundamentally linked to our perception and cognition, with human languages serving as one of the major carriers of world knowledge. In this vein, Large Language Models (LLMs) like ChatGPT epitomize the pre-training of extensive, sequence-based world knowledge into neural networks, facilitating the processing and manipulation of this knowledge in a parametric space. This article explores large models through the lens of “knowledge”. We initially investigate the role of symbolic knowledge such as Knowledge Graphs (KGs) in enhancing LLMs, covering aspects like knowledge-augmented language model, structure-inducing pre-training, knowledgeable prompts, structured CoT, knowledge editing, semantic tools for LLM and knowledgeable AI agents. Subsequently, we examine how LLMs can amplify traditional symbolic knowledge bases, encompassing aspects like using LLM as KG builder and controller, structured knowledge pretraining, LLM-enhanced symbolic reasoning, and the amalgamation of perception with cognition. Considering the intricate nature of human knowledge, we advocate for the creation of Large Knowledge Models (LKM), specifically engineered to manage diversified spectrum of knowledge structures. This ambitious undertaking could entail several key challenges, such as disentangling knowledge representation from language models, restructuring pre-training with structured knowledge, and building large commonsense models, among others. We finally propose a five-“A” principle to distinguish the concept of LKM.
@@ 1.2 Language Models vs Knowledge Graphs
Knowledge Graphs (KGs) model world entities, mental concepts and their relationships in form of graph structures, which are typically extracted or derived from natural language descriptions. With many inspirations from traditional symbolic AI, KGs integrate natural language descriptions and structure knowledge. Typical structural forms include: hierarchical structures(e.g., concept graphs), relational structures(e.g., relational entity graphs), temporal structures(e.g., event logic graphs), etc. Similar to natural languages, KGs represent knowledge in an explicitly symbolic way, thereby offering higher reliability and interpretability to humans.
@@
In this article, we aim to investigate large models through the lens of “knowledge”. Our first focus is on how KGs can enhance LLMs. Topics include knowledge-augmented language modeling, structure-inducing pre-training, knowledgeable prompts, structured Chains of Thought (CoT), semantic tools for LLMs, editing knowledge in large models, and the development of knowledgeable AI agents. We further examine the role of LLMs in enriching traditional symbolic knowledge bases, encompassing the utilizing of LLMs as KG builders and controllers, structured knowledge pre-training, LLM-driven symbolic reasoning, and the integration of perception and cognition. Acknowledging the intricate nature of human knowledge, we then conceptualize a Large Knowledge Model (LKM) in purpose to handle the diversity of structures of human knowledge. This may involve approaches of decoupling knowledge from language models, restructure pretraining with structure knowledge, building large commonsense model, among others. To distinguish the concept of LKM, we propose a five-“A” principle from the aspects of Augmented pretraing, Authentic knowledge, Accountable reasoning, Abundant knowledge coverage, and Aligned with knowledge.
2 Knowledge-Augmented Language Models
@@
The first step towards Large Knowledge Models is to augment LLM with knowledge. In this context, “knowledge” specifically refers to more standardized semantic representations such as domain teminology or ontologies, more structured knowledge representations such as a KG, or richer logic descriptions such as text with CoT. We will discuss how knowledge can be effectively applied in various aspects in the life cycle of a LLM, including pre-training, fine-tuning with instructions, prompt engineering, Chain-of-Thought, AI agents, among others.
2.1 Knowledge Augmentation
@@
Growing numbers of researches have shown that improving the semantic standardization of or adding more structured knowledge into the training corpus would be beneficial to enhance model reasoning capability, either in pre-training or instruction-tuning stages ToG ; MindMap ; InstructProtein ; CoK ; UnifyKGLLM ; ClinGen . For examples, code languages, characterized by their structured nature and abundant computational logic, would be more conducive to activating the reasoning capabilities of large models compared to natural languages PoT ; PoT-CIRS . Similarly, Chain-of-Thought (CoT) prompts CoT ; CoT-SC contain more logical descriptions than ordinary text, thus being more effective in stimulating model reasoning ability. Further research reveals that CoT prompts are more effective when the model scale to a certain extent, which can be explained as that a larger model contains more knowledge for reasoning, and the activation of these knowledge requires more logical prompts.
@@
To delve deeper into the correlations between code structure and model reasoning, we introduce a study called “Program of Thoughts” PoT-CIRS . This research begins by establishing metrics for quantifying the complexity of code structure and code logic. For instance, codes featuring loop operations are considered as having richer computational logic and assigned with higher scores compared to those with mere variable assignment operations. Subsequently, we compile a series of code prompts with different levels of structure or logic scores, and assess their impact on model reasoning performance. Experimental analyses on various benchmarks reveal that the code prompts with higher structure or logic scores are more conducive to improving model reasoning performance.
Code language is not suitable to represent common-sense knowledge, while Knowledge Graphs (KGs) are more adept in modeling such type of knowledge. To further explore the influence of knowledge structure on common-sense reasoning, we transform structured knowledge in a KG into corresponding code constructs in a study called KG2Code CodeKGC . For example, we made classes and properties in a KG transformed to corresponding Class and member variable declarations found in conventional object-oriented programming languages. We then use the code-formed KGs as prompts or directly inject them into base model pre-trainings, and evaluate their impact on downstream tasks. Experimental results have indicated that models augmented with code-formed KGs exhibit notable performance improvements in various dimensions, particularly in extraction tasks.
External knowledge can also enhance the generalization ability of large models. As discussed in RetroPrompt RetroPrompt ,large models usually rely on “mechanical memorization” to handle long-tailed or isolated samples rather than actually learning underlying patterns. This is one of the essential reasons why many large models have poor generalization ability in few-shot scenarios. It’s like a student who only knows how to memorize things mechanically is relatively weak to extrapolate from one instance to others. RetroPrompt addresses this issue by decoupling knowledge from memory, and train a external knowledge base (not necessarily a KG here) independently. Prompts are then enriched by retrieving pertinent knowledge the pretrained knowledge base. Experiments indicate that RetroPrompt manages to effectively improve the generalization ability of models particularly in few-shot scenarios.
2.2 Structure Inducing Pre-training
@@
Structures do matter in pretraining. In a study called “Structure-Inducing Pretraining” StructIndPretrain , the role of imposing relational structure in pretrained language models is investigated. The basic ideas involve crafting connections, either at inter-sample or intra-sample level, to enforce additional structural constraints on the geometric relationships within or between training samples. These constraints range from shallow, implemented via auxiliary pretraining objectives, to more intricate, implicitly integrated through strategies like data augmentation or sample-level contrastive learning.
@@
A typical example of inter-sample structure constraints manifests in the injection of a protein-protein interaction network into the pretraining of a protein language model. This injection infuses the model with sample-level structural insights, i.e. the relational structures of protein interactions. A typical example of intra-sample structure manifests in the linking of entities within a sentence to an external knowledge graph, thereby introducing correlation information in a KG into the training sentences. Both theoretical analysis and experimental verifications are performed and indicate a consistent improvement across a variety of benchmarks.
@@ 2.3.1 Prompt Engineering as Knowledge Engineering
Prompts refer to natural language text describing the instructions that an AI should perform, the context of instructions that an AI can inquire of, or the guidance to think that an AI can follow InstructGPT ; CoT ; FLAN . It has been proven by numerous practices Auto-CoT ; APE ; AutoPrompt ; KNNPrompt that the quality of prompts has a huge impact on model outputs. By providing clear prompts or instructions with specific details, the model can more accurately comprehend human intentions, leading to outputs that are more aligned with human instructions. In contrast to the basic and shallow semantic understanding fostered during base model training, prompts in the fine-tuning stage embed more sophisticated and task-specific knowledge.
@@ 2.3.2 Structured Chain of Thoughts
2.3.3 KG2Instructions
Instruction tuning InstructGPT ; FLAN refers to further training base models with task-specific instructions, typically converted from task-specific training data. Since data from various tasks are converted to an unified input form as instructions or prompts, intruction-tuning with massive downstream task would significantly improve model’s generalization ability, that is, the ability to solve new problems or to complete new tasks.
@@ 2.3.3 KG2Instructions
On the other hand, KGs can be directly converted to instructional training datasets. As exemplified by KG2Instructions proposed in DeepKE-LLM knowlm , a KG refined from WikiData is employed to automatically generate vast amount of training instructions for knowledge extraction. Essentially, these instructions converted from the KG represent a novel form of distant supervision, which leverages the structured information in KGs to generate instructions that guide models in understanding and extracting relevant information across diverse contexts and tasks. Experiments demonstrate notable efficacy in enhancing the generalizability of extraction capabilities of large models.
@@
2.4 Knowledge Editing
A well-known problem of large models is hallucination, TrustworthyLLMs ; KGHallucination , where models generate responses based not on explicit textual evidence rather than on the amalgamation of implicit parameters within the neural network. Once the model goes wrong, it is neither easy to pinpoint the source of errors, nor trivial to correct the errors in a parameterized space. Even more seriously, due to the interconnected nature of knowledge, even minor adjustments to the parameterized knowledge can lead to widespread and unpredictable effects of “a slight move in one part may affect the situation as a whole”. This phenomenon is more pronounced in LLM than in traditional symbolic knowledge bases such as KGs, where changes tend to be more localizable and manageable.

### v2

@@ Abstract
Humankind’s understanding of the world is fundamentally linked to our perception and cognition, with human languages serving as one of the major carriers of world knowledge. In this vein, Large Language Models (LLMs) like ChatGPT epitomize the pre-training of extensive, sequence-based world knowledge into neural networks, facilitating the processing and manipulation of this knowledge in a parametric space. This article explores large models through the lens of “knowledge”. We initially investigate the role of symbolic knowledge such as Knowledge Graphs (KGs) in enhancing LLMs, covering aspects like knowledge-augmented language model, structure-inducing pre-training, knowledgeable prompts, structured CoT, knowledge editing, semantic tools for LLM and knowledgeable AI agents. Subsequently, we examine how LLMs can boost traditional symbolic knowledge bases, encompassing aspects like using LLM as KG builder and controller, structured knowledge pretraining, and LLM-enhanced symbolic reasoning. Considering the intricate nature of human knowledge, we advocate for the creation of Large Knowledge Models (LKM), specifically engineered to manage diversified spectrum of knowledge structures. This promising undertaking would entail several key challenges, such as disentangling knowledge base from language models, cognitive alignment with human knowledge, integration of perception and cognition, and building large commonsense models for interacting with physical world, among others. We finally propose a five-“A” principle to distinguish the concept of LKM.
@@ 1.2 Language Models vs Knowledge Graphs
Knowledge Graphs (KGs) model world entities, mental concepts and their relationships in form of graph structures, which are typically extracted or derived from natural language descriptions. With many inspirations from traditional symbolic AI, KGs integrate natural language descriptions and structure knowledge. Typical structural forms include: hierarchical structures(e.g., concept graphs), relational structures(e.g., relational entity graphs), temporal structures(e.g., event logic graphs), etc. Akin to natural languages, KGs encode knowledge symbolically, ensuring robust reliability, traceable reasoning, and human-centric interpretability.
@@
As depicted in Figure 1, both LLMs and KGs are specifically developed for the representation and manipulation of knowledge. KGs provide highly accurate and interconnected knowledge, enabling controlled reasoning and enhanced human readability. Conversely,LLMs offer expansive coverage of knowledge, exhibit enhanced task generalization capabilities, and utilize a neural representation that optimizes machine efficiency. Yet, KGs face scalability challenges, are limited in their ability to extrapolate, and struggle with reasoning transferability. LLMs, despite their power, incur significant training costs, struggle with deep logical reasoning, and are susceptible to hallucination errors.
Figure 2: An Outline of the Whole Content Structure
@@
In this article, we aim to investigate large models through the lens of “knowledge”. Our first focus is on how KGs can enhance LLMs. Topics include knowledge-augmented language modeling, structure-inducing pre-training, knowledgeable prompts, structured Chains of Thought (CoT), semantic tools for LLMs, editing knowledge in large models, and the development of knowledgeable AI agents. We further examine the role of LLMs in enriching traditional symbolic knowledge bases, encompassing the utilizing of LLMs as KG builders and controllers, structured knowledge pre-training, LLM-driven symbolic reasoning. Acknowledging the intricate nature of human knowledge, we then conceptualize a Large Knowledge Model (LKM) in purpose to handle the diversity of structures of human knowledge. This may involve approaches of decoupling knowledge from language models, restructure pretraining with structure knowledge, building large commonsense model, among others. To distinguish the concept of LKM, we propose a five-“A” principle from the aspects of Augmented pretraing, Authentic knowledge, Accountable reasoning, Abundant knowledge coverage, and Aligned with knowledge.
2 Knowledge-Augmented Large Models
@@
The first step towards Large Knowledge Models is to augment LLM with knowledge. In this context, “knowledge” specifically refers to more standardized semantic representations such as domain terminology or ontologies, more structured knowledge representations such as a KG, or richer logic descriptions such as text with CoT. We will discuss how knowledge can be effectively applied in various aspects in the life cycle of a LLM, including pre-training, fine-tuning with instructions, prompt engineering, Chain-of-Thought, AI agents, among others.
2.1 Expressiveness vs. Scaling Law: The Dichotomy in Knowledge Representation
Traditional research in symbolic Knowledge Representation (KR) mainly focuses on the link between representation expressiveness and reasoning ability, suggesting that more expressive or complex representations facilitate stronger reasoning. However, this perspective often overlooks a vital factor: the scale of knowledge also significantly influences reasoning. For example, the scaling law in large models kaplan2020scaling indicates that diversified reasoning abilities only emerge when the model reaches a certain size. This phenomenon can be attributed to the fact that an increase in model size results in an increase of knowledge coverage, which in turn enables more sophisticated reasoning tasks.
Figure 3: The Dichotomy in Knowledge Representation
Indeed, both expressiveness and scale act as balancing factors for reasoning capabilities as illustrated in Figure 3. More intricate representations result in more precise knowledge, like an ontology for KGs, thus enabling stronger reasoning, but incuring more challenging to acquire the knowledge in large scale. In contrast, simpler representations, like textual sequences for training large models, facilitate extensive coverage of knowledge but compromise delicate reasoning due to reduced precision. As we will delve deeper, this inherent contradiction forms the basis of the complementary nature of KGs and large models, driving advancements in augmenting large model with conventional symbolic KR.
2.2 LLM Pre-training with Knowledge Structure
Let’s first address how to enhance the pre-training stage with various forms of knowledge.
2.2.1 Code Structure and Thought Logic
@@
Numerous studies suggest that enhancing corpus structure or incorporating more logic into the training corpus can improve the reasoning performance of large models ToG ; MindMap ; InstructProtein ; CoK ; UnifyKGLLM ; ClinGen . For examples, code languages, characterized by their structured nature and abundant computational logic, would be more conducive to activating the reasoning capabilities of large models compared to natural languages PoT ; PoT-CIRS . Similarly, Chain-of-Thought (CoT) prompts CoT ; CoT-SC contain more logical descriptions than ordinary text, thus being more effective in stimulating model reasoning ability. Further research reveals that CoT prompts are more effective when the model scale to a certain size, which can be explained as that a larger model contains more knowledge for reasoning, and the activation of these knowledge requires more logical prompts.
@@
To delve deeper into the correlations between code structure and model reasoning, we introduce a study about “Reasoning with Program of Thoughts” PoT-CIRS . This research begins by establishing metrics for quantifying the complexity of code structure and code logic. For instance, codes featuring loop operations are considered as having richer computational logic and assigned with higher scores compared to those with mere variable assignment operations. Subsequently, we compile a series of code prompts with different levels of structure or logic scores, and assess their impact on model reasoning performance. Experimental analyses on various benchmarks reveal that the code prompts with higher structure or logic scores are more conducive to improving model reasoning performance.
Although the representation of code structure and CoT logic remains at a relatively simple level, it has nonetheless contributed to enhancing the reasoning capabilities of large models. It can be anticipated that the utilization of data representations with deeper logic and more complex structure, such as knowledge graphs, can further enhance the capabilities of models.
2.2.2 Knowledge-injecting Pre-training
Large models build their reasoning prowess by understanding sequences of words. To achieve deep reasoning similar to human cognitive processes, it’s essential to progress beyond basic word semantics to the formation and comprehension of advanced concepts, entities, and their interconnections. Injecting knowledge graph with pre-training models is a promising strategy to enable such type of richer understanding and move beyond the limitations of simple word co-occurrences to embrace the complexity of higher-order reasoning. There exist several strategies for integrating knowledge graphs into pre-trained language models, broadly categorized based on the stage of integration: input layer, architecture layer, and objective layer as illustrated in Figure 4.
Figure 4: Knowledge Injection on Different Layers.
  • Input-layer Injection. At the input layer, the simplest strategy is to convert knowledge graphs into natural language descriptions, thus expanding the model’s input without altering its architecture or training process. Notable examples include K-BERT kbert and CoLAKE CoLAKE , among others. The core idea revolves around leveraging the relationships between entities in knowledge graphs to enrich the contextual information of entities within sentences. Typically, the process begins with establishing connections between entities in sentences and those in the knowledge graph using entity linking technology. Subsequently, sentences are expanded with triples corresponding to these knowledge graph entities before being fed into the large model. Injecting knowledge at the input layer fundamentally serves as a form of data augmentation. These methods do not require modifications to the model’s architecture and are compatible with various off-the-shelf models, making them easy to implement. Additionally, because both structured and textual knowledge are trained within the same model, this approach avoids mismatches between the textual representation space and the structured knowledge representation space. However, converting inherently structured graphs into text sequences inevitably leads to some loss of semantic representation, and the utilization of structural signals is not fully realized.
  • Intermediate-layer Injection. The second strategy involves initially employing structured knowledge representation learning methods, such as KG Embedding, to obtain vector representations of entities and relationships, which are then integrated into language models. Notable examples include ERNIE ERNIE , KnowBERT KnowBERT , KG-BART KG-BART , KT-NET KT-NET , and BERT-MK Bert-MK . For instance, ERNIE starts by pre-training entity representations using TransE. The original textual input remains unchanged, following alignment and interaction with these entity vectors at the intermediate layer of the model with a simple attention mechanism. Since the entity vector representations have already learned the structured features from the original knowledge graphs, this approach better preserves the graph’s structural semantics. However, since textual semantics and graph structure representations are derived from two different pre-training models, there might be interference between textual semantic signals and knowledge structure signals, potentially acting as noise to each other.
  • Objective-layer Injection. The third strategy employs knowledge injection at the objective level, often through multi-task learning and refined LOSS optimization. Notable examples include KEPLERKEPLER , WKLM WKLM , JAKET JAKET , and BERT-MK Bert-MK . For instance, KEPLER adds a KG embedding Loss on top of the standard masked prediction Loss, using a multi-task learning framework for simultaneous training. As the same encoder is used for both text and entity information, and a single Loss governs the learning of both KG embedding and language model pre-training, this method preserves more structural signals compared to input layer injection and avoid the need for explicit alignment between text and KG required by architecture layer injection.
2.2.3 Structure-inducing Pre-training
@@
We have already mentioned that human knowledge is not a simple sequence of texts, but rather has a rich structure. In a study called “Structure-Inducing Pretraining” StructIndPretrain , the role of imposing relational structure in pretrained language models is investigated. The basic ideas involve crafting connections, either at inter-sample or intra-sample level, to enforce additional structural constraints on the geometric relationships within or between training samples. These constraints range from shallow, implemented via auxiliary pretraining objectives, to more intricate, implicitly integrated through strategies like data augmentation or sample-level contrastive learning.
@@
A typical example of inter-sample structure constraints manifests in the injection of a protein-protein interaction network into the pretraining of a protein language model. This injection infuses the model with sample-level structural insights, i.e. the relational structures of protein interactions. A typical example of intra-sample structure manifests in the linking of entities within a sentence to an external knowledge graph, thereby introducing correlation information in a KG into the training sentences. Both theoretical analysis and experimental verifications are performed and indicate a consistent improvement with structure-inducing across a variety of benchmarks.
Figure 5: Structure-Augmented Pre-training
2.2.4 reStructured Pre-training
Another relevant research is so-called reStructured Pre-training yuan2022restructured . Rather than imposing external structural knowledge, as in those structure-inducing methods, reStructure approach delves into the structural signals inherent in the data itself. The world’s data is rich with various structural signals in different formats. For instance, it could be as straightforward as the “next token” signal in a sentence, metadata for a given text, factual knowledge in a triple format, entity or relation annotations, dependency structures between words, and numerous others. To make LLMs better educated, it is justifiable to mine these structure signals automatically or semi-automatically, and restructure all of them into unified forms for pre-training models.
@@ 2.3 Knowledgeable Prompts
Next, we will explore how to utilize knowledge to enhance prompt instructions.
@@ 2.3.1 Prompt Engineering as Knowledge Engineering
Prompts refer to natural language text describing the instructions that an AI should perform, the context of instructions that an AI can inquire of, or the guidance to think that an AI can follow InstructGPT ; CoT ; FLAN . It has been proven by numerous practices Auto-CoT ; APE ; AutoPrompt ; KNNPrompt that the quality of prompts has a huge impact on model outputs. By providing prompts or instructions with more specific details, the model can more accurately comprehend human intentions, leading to outputs that are more aligned with human instructions.
@@ 2.3.2 Structured Chain of Thoughts
2.3.3 KG to Instructions
Instruction tuning InstructGPT ; FLAN refers to further training base models with task-specific instructions, typically converted from task-specific training data. Since data from various tasks are converted to an unified input form as instructions or prompts, instruction-tuning with massive downstream task would significantly improve model’s generalization ability, that is, the ability to solve new problems or to complete new tasks.
@@ 2.3.3 KG to Instructions
On the other hand, KGs can be directly converted to instructional datasets. As exemplified by DeepKE-LLM knowlm , a KG refined from WikiData is employed to automatically generate vast amount of training instructions for knowledge extraction. Essentially, these instructions converted from the KG represent a novel form of distant supervision, which leverages the structured information in KGs to generate instructions that guide models in understanding and extracting relevant information across diverse contexts. As shown in Figure 6, InstructProtein wang2023instructprotein introduces a knowledge graph-based instruction generation framework to construct a high-quality instruction datasets for training a protein language model . In particular, the instructions inherit the structural relations between proteins and function annotations in a protein knowledge graphs, which empowers the model to engage in the causal modeling of protein functions, akin to the chain-of-thought processes in natural languages.
Figure 6: InstructProtein: an example of building instruction dataset with a KG InstructProtein
@@
2.4 KG-RAG: Knowledge-based Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) refers to the enhancement of large models’ generative capabilities through querying an external knowledge base, which could be external search engines, domain-specific corpora, structured databases, or knowledge graphs. Due to its cost-effectiveness in training, avoidance of model hallucinations, and compliance with privacy and copyright standards, RAG has become an essential tool in large model applications.
Knowledge-augmented RAG leverages the conceptual hierarchies, entity connections, and logical rules within knowledge graphs to improve retrieval processes. The essence of RAG involves segmenting documents or images into data chunks, which are embedded and stored in a vector database for retrieval. The primary issue with chunking lies in the risk of losing the document’s inherent logic and contextual relationships. Vector representations of these chunks offer merely a basic feature vector for similarity computation, falling short of capturing the deeper semantic connections among entities or concepts—a gap knowledge graphs were designed to bridge.
A practical way to apply KG-RAG is through the creation of entity or concept trees, enhancing RAG into versions like Entity-RAG or Topic-RAG. Using knowledge graphs to build hierarchical structures of concepts, topics, or entities, and connecting these elements within document chunks to the knowledge graphs, can significantly improve the retrieval process through these organized knowledge hierarchies. A typical example is Tree-RAG Tree-RAG , which merges RAG search with an entity tree that enhances contextual retrieval in tandem with a vector database. This entity tree catalogs the hierarchy of organizational structures and their categories, such as departments and their subdivisions. The entity tree not only clarifies an entity’s context within an organization but also boosts search across document chunks at an entity level, thus deepening the model’s comprehension of entity relationships.
Another KG-RAG approach uses an entity graph to augment RAG. Initially, document chunks are processed using knowledge extraction to create a temporary, document-specific knowledge graph for RAG enhancement. When an external knowledge graph like WikiData is available, it’s merged with the local graph and stored into a graph database. During retrieval, queries access both vector and graph databases to find relevant document chunks and associated triples, providing context for the large model to generate responses. KG-FiD KG-FiD exemplifies this by blending retrieved paragraphs with external knowledge graphs. It firstly create a comprehensive KG that contains both paragraph-specific and external knowledge. Graph Neural Networks (GNN) are then used to semantically analyze and iteratively re-rank paragraphs based on their relevance, filtering out irrelevant content. This filtered content is then fed into the decoder as context, significantly improving response accuracy.
External knowledge can also enhance the generalization ability of large models. As discussed in RetroPrompt RetroPrompt ,large models usually rely on “mechanical memorization” to handle long-tailed or isolated samples rather than actually learning underlying patterns. This is one of the essential reasons why many large models have poor generalization ability in few-shot scenarios. It’s like a student who only knows how to memorize things mechanically is relatively weak to extrapolate from one instance to others. RetroPrompt addresses this issue by decoupling knowledge from memory, and train a external knowledge base (not necessarily a KG here) independently. Prompts are then enriched by retrieving pertinent knowledge from the pretrained knowledge base. Experiments indicate that RetroPrompt manages to effectively improve the generalization ability of models particularly in few-shot scenarios.
2.5 KGs as References for Hallucination Mitigation
A well-known problem of large models is hallucination TrustworthyLLMs ; KGHallucination , where models generate responses based not on explicit textual evidence rather than on the amalgamation of implicit parameters within the neural network. Tackling these issues requires detecting those hallucination first. Knowledge graphs, known for their high accuracy and clear logical definitions due to manual verification, serve as an effective reference for detecting such issues. Their utility spans three key areas:
  • Fact Checking: This step compares the model’s responses with the knowledge graph’s triples to verify accuracy, relying on the graph’s comprehensive coverage.
  • Logical Consistency: Utilizing the knowledge graph’s logical structures to enhances the capacity to identify inaccuracies by examining the logical consistency of responses.
  • Validation Model Training: The strong semantic and logical integrity can be utilized to train a discriminator model that can assess or judge the accuracy of the generated content.
Knowledge graphs thus play a crucial role in improving large models’ reliability by identifying hallucinations and detecting inaccuracies efficiently.
2.6 Large Model Knowledge Editing
In addition to detecting hallucinations, addressing this issue extends to directly correcting them. Just as we could query and interact with knowledge graphs in a symbolic space, we would also like to interrogate and manipulate the knowledge stored in the parameterized space of large models through query, edit, modify, delete, or update it.